#2 The following items were shown on the balance sheet of Easton Corporation on December 31, 2012: Stockholders equity Paid-in capital Capital stock Common stock, $5 par value, 400,000 shares Authorized; . shares issued and outstanding Additional paid-in capital In excess of pa Total paid-in capitail. Total paid-in capital and retained earnings Total stockholders equity 2,015,000 Retained earnings.... 750,000 . 2,765,000 Less: Treasury stock (18,000 shares... (180,000) Instructions Complete the following statements and show your computations. (a) The number of shares of common stock issued was
media%2Fa57%2Fa577913c-6ca4-43b4-9f8c-8d
Answer #1

(a) The number of shares of common stock issued was 370,000.

Common stock/Par value per share = $1,850,000/$5 = 370,000

(b) The number of shares of common stock outstanding was 352,000.

Common stock issued – Treasury stock = 370,000 – 18,000 = 352,000

(c) The sales price of the common stock when issued was $2,015,000

Common stock + Additional paid-in capital = $1,850,000 + $165,000 = $2,015,000

(d) The cost per share of the treasury stock was $10.

Treasury stock $/Number of treasury stock = $180,000/18,000 = $10

(e) The average issue price of the common stock was $5.45

Total paid-in capital/Number of shares issued = $2,015,000/370,000 = $5.45

(f) Assuming that 25% of the treasury stock is sold at $20 per share, the balance in the Treasury Stock account would be $135,000.

Treasury stock sold = 25% x 18,000 shares = 4,500

Cost of treasury stock sold = 4500 x $10 = $45,000

Balance in Treasury stock account = $180,000 - $45,000 = $135,000